# Language Model Planning from an Information Theoretic Perspective

### v1

@@ Abstract
The extent to which decoder-only language models (LMs) engage in planning, that is, organizing intermediate computations to support coherent long-range generation, remains an open and important question, with implications for interpretability, reliability, and principled model design. Planning involves structuring computations over long horizons, considering multiple possible continuations, and selectively reusing past information, but how effectively transformer-based LMs realize these capabilities is still unclear. We address these questions by analyzing the hidden states at the core of transformer computations, which capture intermediate results and act as carriers of information. Since these hidden representations are often redundant and encumbered with fine-grained details, we develop a pipeline based on vector-quantized variational autoencoders that compresses them into compact summary codes. These codes enable measuring mutual information, allowing systematic analysis of the computational structure underlying model behavior. Using this framework, we study planning in LMs across synthetic grammar, path-finding tasks, and natural language datasets, focusing on three key aspects: (i) the planning horizon of pre-output computations, (ii) the extent to which the model considers alternative valid continuations, and (iii) the reliance of new predictions on earlier computations. By answering these questions, we advance the understanding of how planning is realized in LMs and contribute a general-purpose pipeline for probing the internal dynamics of LMs and deep learning systems. Our results reveal that the effective planning horizon is task-dependent, that models implicitly preserve information about unused correct continuations, and that predictions draw most on recent computations, though earlier blocks remain informative.
@@ 1 Introduction
Language models (LMs) have advanced so rapidly that they now engage in open-ended conversation, write functional code, and even solve challenging mathematical problems [9, 1, 20, 55]. Beyond these raw capabilities, scaffolding and augmentation techniques further enhance reasoning and planning [64, 63, 46, 65, 7]. In parallel, hybrid systems integrating LMs with symbolic planners or external tools have achieved state-of-the-art performance in embodied and tool-use domains [67, 60, 48].
All the above success suggests that LMs have a certain capability of “planning ahead,” similar to what humans do in generating coherent long speeches and long-horizon problem solving. Yet, the typical training method for LMs, next token prediction, focuses only on predicting the next token, which seemingly suggests LMs are myopic. Motivated by this contrast, this paper seeks to understand the planning behavior of LMs and how it is affected by next-token prediction (NTP) vs. multi-token prediction (MTP) in training. As “planning behavior” is a broad term that has many facets, we further narrow our focus to the following three canonical aspects of planning. First, good planners are forward-looking. They structure intermediate computations over a receding horizon so that near-term actions serve longer-term objectives, an idea made explicit in model-predictive control (MPC) and world-model–based approaches [38, 35, 22, 28]. Second, good planners are branch-aware. Before committing, they keep multiple plausible futures “alive,” comparing candidates rather than greedily following a single line, an ability central in classical MPC and world-model–based control methods [38, 35, 23], and closely tied to robustness under paraphrase and logically related prompts [32, 2, 45] and to the success of search-based scaffolds such as Tree- and Graph-of-Thoughts [65, 7]. Third, good planners are stateful. They reuse and build upon earlier computations so that later decisions depend coherently on what has already been processed. In classical planning and control, this corresponds to carrying forward simulated states, costs, or latent variables across the planning horizon so that each new action is consistent with past predictions [38, 35]. Given the above three aspects about planning, this paper seeks to understand,
To what extent are LMs forward-looking, branch-aware, and stateful?
@@
We approach the above questions by studying the internal states and computations of an LM, as recent work shows that transformers internally encode rich, high-level abstractions such as belief states, game configurations, and world models that extend far beyond their immediate outputs, and that these abstractions can be partially extracted [47, 29, 41, 44]. This makes the internal representations of LMs a natural locus for investigating how they plan and reason. While not aimed at these specific questions, existing methods have explored internal computations of models, e.g. using probing classifiers to test for linguistic features in hidden states or mechanistic analyses to identify circuits and disentangled features within transformers [25, 15, 8]. However, these only indirectly relate to our questions, and despite important progress, these strategies have clear limitations. Circuit discovery requires heavy manual engineering [15, 61, 10, 36], while probing risks conflating genuine representations with probe artifacts [25, 58, 42, 27, 26]. These limitations necessitate new approaches towards understanding the internal computations that are automated, scalable, and less susceptible to probing’s confounding of representations learned by the probe with those in the model itself.
@@
Contribution 1: An information-theoretic tool towards understanding internal states of LMs: In contrast to prior works in mechanistic interpretability or probing methods, we propose a framework to study how planning-related computations are organized inside LMs that is both automated (no manual circuit engineering) and free from confounding caused by learned probes. Specifically, (a) to avoid confounding caused by the probe, we propose to calculate the Mutual Information (MI) between two components of the LMs, which provides a lens on how planning-related information is organized. For example, to understand the “forward-looking” question, the MI between the internal states of the prefix and that of future generations can shed light on how much computations inside LMs plan ahead for future tokens. In contrast, probing detects whether a piece of information is present in a LM hidden state by training an external model to predict the information from the hidden state, which may introduce additional representational power brought by the external model. On the other hand, the MI is a confound-resistant metric measuring how much two variables contain information about each other, which avoids confounding by not introducing a new model. Moreover, unlike correlation, which is sensitive primarily to linear relationships under a chosen parameterization, MI captures arbitrary statistical dependence and is symmetric and invariant under invertible reparameterizations. (b) To make the MI calculation scalable to potentially very high dimensional hidden state vectors, we employ a Vector-Quantized Variational Autoencoder (VQ-VAE) to map collections of block outputs into discrete codes that serve as coarse summaries of internal states [57]. We use VQ-VAE as a practical and suitable compressor with a discrete codebook and transformer encoder, enabling MI estimation over variable-sized blocks; our validation study (App. A.3) supports this choice. We calculate MI between the coarse summaries instead of the raw hidden states. This step is crucial: fine-grained activations are high-dimensional and redundant, making a detailed direct analysis infeasible, while the compressed codes capture the salient distinctions necessary for comparing computations across layers, positions, and contexts.
While our framework is general and could be applied to other applications in deep learning, such as disentangling reasoning circuits in multimodal architectures or characterizing information flow in reinforcement learning (RL), in this work, we use it to analyze the planning of transformer models, which we introduce below.
@@
Contribution 2: Understanding the planning of LMs. Using the information-theoretic framework, we analyze LMs’ ability to be (i) forward-looking, (ii) branch-aware, and (iii) stateful in decoder-only transformers, across a symbolic task, a structured reasoning problem, and natural text (§ 3):
@@
  • History in the Plan (§ 3.3): to what extent the token generation relies on earlier computations.
Our results show that planning is task-contingent and weakly modulated by the training objective loss, i.e., MTP versus NTP. When the task demands a longer horizon, the LM’s pre-output states retain information about tokens beyond the immediate next step. In contrast, in locally syntactic settings, this dependence concentrates near the next token. Training with MTP loss modestly reduces purely myopic behavior. Internally, the model encodes alternatives to the produced answer, and the strength of this branching awareness of the plan varies with task difficulty and model quality. Finally, next-token decisions draw most strongly on the last layers and the most recent token indices, yet earlier blocks retain nontrivial information, indicating stateful computation.
@@
Despite the recent advances in LM planning abilities [63, 46, 64, 65, 7, 67, 60, 48], recent studies highlight that significant challenges remain, and current approaches to LM planning still fall short of fully addressing complex reasoning and decision-making tasks. For example, [32] show that models can produce conflicting answers under logically related prompts despite local plausibility, which shows lack of long-horizon planning although short-horizon it works fine; [2, 45, 37, 62] highlight inconsistencies in the model’s outputs and struggles with planning tasks. Taken together, these findings underscore that understanding whether and how planning arises in LMs is not only an open empirical challenge, but also central to both their interpretability and the principled design of future model architectures.
@@
To interpret LMs, some approaches treat the model as a black box and design tasks or benchmarks to gauge reasoning, robustness, or generalization abilities at a behavioral level [51, 30]. While such evaluations provide useful insights, they miss several perspectives that can be gained by examining the internal mechanisms of the model. In contrast, mechanistic interpretability seeks to reverse-engineer transformer computations into human-understandable parts, treating the residual stream as the main information pathway and attention heads as separate components that pass information along [15, 40, 12, 8, 25, 14, 33]. However, circuit discovery requires significant manual engineering [61, 10, 36]. Beyond empirical tools, mathematically grounded perspectives also shed light on transformer and LLM interpretability [34, 3, 56, 16], yet these approaches are often criticized for lacking a one-to-one correspondence with experimental results, since their proofs typically rely on strong assumptions that may not hold for the highly non-smooth LM architectures. In contrast to them, our method (§ 2) enables the study of LMs both behaviorally and structurally, by examining their internal mechanisms, without the need for labor-intensive circuit discovery or strong assumptions on the LM. Still, due to its reliance on information-theoretic tools, our approach can only capture aggregate phenomena, providing insights in an average sense rather than interpreting individual input prompts.
@@
A different line of interpretability work views LM hidden states as structured representations that can be "probed"-probing refers to training lightweight models, often linear classifiers, to read out specific information from hidden states in order to test what the model represents internally. Using probing, researchers have shown that transformer hidden states encode structured belief-state and world-model–like information [47, 21, 24]. Other works demonstrate that a single hidden state can carry information about multiple future tokens [41] and that probing can reveal the underlying algorithms LLMs use to solve tasks [4]. However, standard accuracy-based probing has been criticized for combining what the probe can learn with what the representation actually encodes, making the results sensitive to probe capacity, data size, and hyperparameters [25, 42]. Further, high probe scores often come from exploiting superficial linear context cues rather than genuine structural knowledge [27]: probing can even reveal features that a model does not use for its task [43, 26]. Consequently, some critiques motivate adopting information-theoretic lenses [58, 13]. Similar to some of these information-theoretic approaches [54, 49, 50, 59], our method makes use of information theory. For a more detailed discussion of related work, please refer to App. C.
@@ 2 Method
Given a prefix sequence of length $T$, ${\mathbf{x}}_{1:T}=\left[x_{1},x_{2},\dots,x_{T}\right]$, the decoder-only LM, $\mathcal{M}$, generates subsequent tokens $\{\widehat{x}_{T+\tau}\mid\tau=1,2,\dots\}$ autoregressively. We denote the hidden activation at $t$ after the $\ell^{\text{th}}$ transformer layer by ${h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}}\in\mathbb{R}^{d}$ where $d$ is the hidden dimension of the model.
In our analysis, we treat the outputs $h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}$} of individual transformer blocks as the fundamental units of computation [^1] We use the terms “hidden state” and “(hidden) activation” interchangeably to refer to these outputs.. We posit that a block’s output is informative about the computation performed within that block [41, 47, 29, 15, 33]. We use them as a proxy for the model’s internal processing since each block’s computation contributes to the overall function of the model. To analyze how planning emerges in the model’s internal dynamics, we examine collections of layer–token indices $\mathcal{S}$ and aggregate the corresponding block outputs as $G_{\mathcal{S}}=\{{h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}}\mid(\ell,t)\in\mathcal{S}\}$, a representation for the computations within its block. While $G_{\mathcal{S}}$ provides a fine-grained representation of block computations, it often contains unnecessary details for our planning analysis. Hidden states may encode small variations due to token positions, normalization shifts, or attention noise, factors that are critical for exact output reconstruction but irrelevant for studying planning structure. Moreover, MI estimates between such high-dimensional vectors have high variance and are hard to find. To address both issues, we construct coarse, discrete representations of $G_{\mathcal{S}}$. These abstractions preserve the structural dependencies required for our analysis while filtering out irrelevant micro-level variability, yielding more stable and interpretable MI estimates.
@@
To investigate how planning is realized within the model, we use MI as a principled way to quantify relationships between the coarse representations of different computations. Suppose we focus on two hidden state blocks, $G_{{A}}$ and $G_{{B}}$, which are random variables depending on the data distribution and the trained model parameters. We denote their coarse, discrete representations by $Z_{A}\in\mathcal{Z_{A}}$ and $Z_{B}\in\mathcal{Z_{B}}$ where $\mathcal{Z_{A}}$ and $\mathcal{Z_{B}}$ are the finite alphabets of possible codes. For realizations $z_{a}\in\mathcal{Z}_{A}$ and $z_{b}\in\mathcal{Z}_{B}$, $p(z_{a},z_{b})$, $p(z_{a})$, and $p(z_{b})$ refer to the joint probability of $Z_{A}$ and $Z_{B}$, and the marginal probabilities of $Z_{A}$ and $Z_{B}$, respectively. Mutual information, which measures how much knowing $Z_{A}$ reduces uncertainty about $Z_{B}$ (and vice versa), is defined as,
@@
Since the true data distribution is inaccessible, we approximate by replacing $p(z_{a})$, $p(z_{b})$, and $p(z_{a},z_{b})$ in Eq. 1 with their empirical estimates from the datasets. We further define normalized mutual information (nMI) metric as
@@
where $\mathcal{I}_{\max}$ denotes the maximum value across the set of MI calculations performed in the same experimental analysis for comparison. For our purpose of LM planning investigation, the nMI metric is sufficient since our analyses depend on relative, rather than absolute, comparisons between sets of computations. Absolute values can be misleading for three reasons: (i) the estimated MI of coarse representations is bounded above by the logarithm of the alphabet size, making it inherently dependent on the experimental choice; (ii) any coarse-graining necessarily discards some information, which is advantageous for our setting as discussed but removes any meaning of absolute MI values related to the original hidden states; and (iii) answering our three research questions (§ 3) requires only relative comparisons of MI instead of relying on absolute values.

### v2

# Internal Planning in Language Models: Characterizing Horizon and Branch Awareness

@@ Abstract
The extent to which decoder-only language models (LMs) engage in planning, that is, organizing intermediate computations to support coherent long-range generation, remains an important question, with implications for interpretability, reliability, and principled model design. Planning involves structuring computations over long horizons, and considering multiple possible continuations, but how far transformer-based LMs exhibit them without external scaffolds, e.g., chain-of-thought prompting, is unclear. We address these questions by analyzing the hidden states at the core of transformer computations, which capture intermediate results and act as carriers of information. Since these hidden representations are redundant and encumbered with fine-grained details, we develop a pipeline based on vector-quantized variational autoencoders that compresses them into compact summary codes. These codes enable measuring mutual information and analyzing the computational structure of the underlying model behavior. Using this framework, we study planning in LMs across synthetic grammar, path-finding tasks, and natural language datasets, focusing on two planning properties: (i) the planning horizon of pre-output computations, and (ii) the extent to which the model considers alternative valid continuations. As a separate downstream use of the same pipeline, we also analyze how decision-relevant information is distributed across layers and earlier prefix blocks when producing next-token predictions. Together, these analyses advance our understanding of planning in LMs and provide a general-purpose pipeline for inspecting internal model dynamics. Our results reveal that the effective planning horizon is task-dependent, that models implicitly preserve information about unused correct continuations, and that predictions draw most on recent computations, though earlier blocks remain informative.
@@ 1 Introduction
Language models (LMs) have advanced so rapidly that they now engage in open-ended conversation, write functional code, and even solve challenging mathematical problems (Brown, 2020; Achiam et al., 2023; Grattafiori et al., 2024; Touvron et al., 2023). Beyond these raw capabilities, scaffolding and augmentation techniques further enhance reasoning and planning (Wei et al., 2022; Wang et al., 2024c; Sel et al., 2025; Yao et al., 2023; Besta et al., 2024). In parallel, hybrid systems integrating LMs with symbolic planners or external tools have achieved state-of-the-art performance in embodied and tool-use domains (Zhao et al., 2023; Wang et al., 2024a; Shen et al., 2023).
All the above success suggests that LMs have a certain capability of “planning ahead,” similar to what humans do in generating coherent long speeches and long-horizon problem solving. Yet, the typical training method for LMs, next token prediction, focuses only on predicting the next token, which seemingly suggests LMs are myopic. Motivated by this contrast, this paper seeks to understand the planning behavior of LMs and how it is affected by next-token prediction (NTP) vs. multi-token prediction (MTP) in training. As “planning behavior” is a broad term that has many facets, we further narrow our focus to the following two canonical aspects of planning. First, good planners are forward-looking. They structure intermediate computations over a receding horizon so that near-term actions serve longer-term objectives, an idea made explicit in model-predictive control (MPC) and world-model–based approaches (Morari and Lee, 1999; Mayne et al., 2000; Ha and Schmidhuber, 2018; LeCun, 2022). Second, good planners are branch-aware, before committing, they keep multiple plausible futures “alive,” comparing candidates rather than greedily following a single line, an ability central in classical MPC and world-model–based control methods (Morari and Lee, 1999; Mayne et al., 2000; Hafner et al., 2019), and closely tied to robustness under paraphrase and logically related prompts (Lin et al., 2025; Ahn and Yin, 2025; Saxena et al., 2024) and to the success of search-based scaffolds such as Tree- and Graph-of-Thoughts (Yao et al., 2023; Besta et al., 2024). Given the above aspects about planning, this paper seeks to understand,
To what extent are LMs forward-looking and branch-aware?
@@
We approach the above questions by studying the internal states and computations of an LM, as recent work shows that transformers internally encode rich, high-level abstractions such as belief states, game configurations, and world models that extend far beyond their immediate outputs, and that these abstractions can be partially extracted (Shai et al., 2024; Li et al., 2023; Pal et al., 2023; Richens et al., 2025). This makes the internal representations of LMs a natural locus for investigating how they plan and reason. While not aimed at these specific questions, existing methods have explored internal computations of models, e.g. using probing classifiers to test for linguistic features in hidden states or mechanistic analyses to identify circuits and disentangled features within transformers (Hewitt and Liang, 2019; Elhage et al., 2021; Bricken et al., 2023). However, these only indirectly relate to our questions, and despite important progress, these strategies have clear limitations. Circuit discovery requires heavy manual engineering (Elhage et al., 2021; Wang et al., 2023; Chan et al., 2022; Meng et al., 2022), while probing risks conflating genuine representations with probe artifacts (Hewitt and Liang, 2019; Voita and Titov, 2020; Pimentel et al., 2020; Kunz and Kuhlmann, 2020; Kumar et al., 2022). These limitations necessitate new approaches towards understanding the internal computations that are automated, scalable, and less susceptible to probing’s confounding of representations learned by the probe with those in the model itself.
@@
Contribution 1: We propose an information-theoretic framework to study how planning-related computations are organized inside LMs that is both automated (no manual circuit engineering) and free from confounding caused by learned probes. Specifically, to avoid probe-induced confounding, we compute mutual information (MI) between learned discrete representations that summarize internal states of the LM. By comparing these MI relationships, we characterize whether internal computations exhibit patterns consistent with forward-looking and branch-aware computation. For example, to understand the “forward-looking” characteristic, the MI between the internal states of the prefix and that of future generations can shed light on how much computations inside LMs plan ahead for future tokens. In contrast, probing aims to detect whether a piece of information is present in a LM hidden state by training an external model to predict the information from the hidden state, which may introduce additional representational power brought by the external model. The resulting prediction losses are often treated as informal proxies for “how much information” a hidden state contains but they depend on the marginal complexity and scale of the chosen target and can violate basic information-theoretic desiderata such as data processing. As we show in our experiments (App. D.2), such probe-based quantities can be strongly influenced by these nuisance factors and need not track true mutual information even in simple controlled settings. In contrast, computing MI between learned representations gives a confound-resistant metric of how much two variables share information, without introducing an additional supervised model whose capacity or objective may obscure the underlying LM computation. Moreover, unlike correlation, which is sensitive primarily to linear relationships under a chosen parameterization, MI captures arbitrary statistical dependence and is symmetric and invariant under invertible reparameterizations. To make the MI calculation scalable to potentially very high dimensional hidden state vectors, we employ a Vector-Quantized Variational Autoencoder (VQ-VAE) to map collections of block outputs into discrete codes that serve as coarse summaries of internal states (Van Den Oord et al., 2017). We use VQ-VAE as a practical compressor with a discrete codebook and transformer encoder, enabling MI estimation over variable-sized blocks; our validation study (App. A.4) supports this choice. We calculate MI between the coarse summaries instead of the raw hidden states. This step is crucial: fine-grained activations are high-dimensional and redundant, making a detailed direct analysis infeasible, while the compressed codes capture the salient distinctions necessary for comparing computations across layers, positions, and contexts. Beyond these planning analyses, the same pipeline also supports a practical diagnostic: localizing where next-token decision information resides across layers and earlier prefix blocks. While our framework is general and could be applied to other applications in deep learning, in this work, we use it to analyze planning in transformer models, and we additionally demonstrate the localization diagnostic as a method application.
@@
Contribution 2: Understanding the planning of LMs. Using the information-theoretic framework, we analyze LMs’ ability to be (i) forward-looking, (ii) branch-aware, and (iii) as an application of our framework, how layers and earlier prefix blocks contain next token information, across a symbolic task, a structured reasoning problem, and natural text (§ 3):
@@
  • Diagnostic of the information in the computational history (§ 3.3): where decision-relevant information about the next token is concentrated across layers and earlier prefix blocks.
Our results show that planning is task-contingent and weakly modulated by the training objective loss, i.e., MTP versus NTP. When the task demands a longer horizon, the LM’s pre-output states retain information about tokens beyond the immediate next step. In contrast, in locally syntactic settings, this dependence concentrates near the next token. Training with MTP loss modestly reduces purely myopic behavior. Internally, the model encodes alternatives to the produced answer, and the strength of this branching awareness of the plan varies with task difficulty and model quality. Finally, next-token decisions draw most strongly on the last layers and the most recent token indices.
@@
Despite the recent advances in LM planning abilities (Wang et al., 2024c; Sel et al., 2025; Wei et al., 2022; Yao et al., 2023; Besta et al., 2024; Zhao et al., 2023; Wang et al., 2024a; Shen et al., 2023), recent studies highlight that significant challenges remain, and current approaches to LM planning still fall short of fully addressing complex reasoning and decision-making tasks. For example, Lin et al. (2025) show that models can produce conflicting answers under logically related prompts despite local plausibility, which shows lack of long-horizon planning; Ahn and Yin (2025); Saxena et al. (2024); Momennejad et al. (2023); Wang et al. (2024b) highlight inconsistencies in the model’s outputs and struggles with planning tasks. Taken together, these findings underscore that understanding whether and how planning arises in LMs is not only an open empirical challenge, but also central to both their interpretability and the principled design of future model architectures.
@@
To interpret LMs, some approaches treat the model as a black box and design tasks or benchmarks to gauge reasoning, robustness, or generalization abilities at a behavioral level (Srivastava et al., 2023; Liang et al., 2022). While such evaluations provide useful insights, they miss several perspectives that can be gained by examining the internal mechanisms of the model. In contrast, mechanistic interpretability seeks to reverse-engineer transformer computations into human-understandable parts, treating the residual stream as the main information pathway and attention heads as separate components that pass information along (Elhage et al., 2021; Olsson, 2022; Cunningham et al., 2023; Bricken et al., 2023; Hewitt and Liang, 2019; Dunefsky et al., 2024; Lindsey et al., 2025). However, circuit discovery requires significant manual engineering (Wang et al., 2023; Chan et al., 2022; Meng et al., 2022). Beyond empirical tools, mathematically grounded perspectives also shed light on transformer and LLM interpretability (Liu et al., 2023; Ahn et al., 2023; Ustaomeroglu and Qu, 2025; Gao et al., 2024), yet these approaches are often criticized for lacking a one-to-one correspondence with experimental results, since their proofs typically rely on strong assumptions that may not hold for the highly non-smooth LM architectures. In contrast to them, our method (§ 2) enables the study of LMs both behaviorally and structurally, by examining their internal mechanisms, without the need for labor-intensive circuit discovery or strong assumptions on the LM. Still, due to its reliance on information-theoretic tools, our approach can only capture aggregate phenomena, providing insights in an average sense rather than interpreting individual input prompts.
@@
A different line of interpretability work views LM hidden states as structured representations that can be ”probed”-probing refers to training lightweight models, often linear classifiers, to read out specific information from hidden states in order to test what the model represents internally. Using probing, researchers have shown that transformer hidden states encode structured belief-state and world-model–like information (Shai et al., 2024; Gurnee and Tegmark, 2024; Hazineh et al., 2023). Other works demonstrate that a single hidden state can carry information about multiple future tokens (Pal et al., 2023; Wu et al., ) and that probing can reveal the underlying algorithms LLMs use to solve tasks (Allen-Zhu, 2024). Although there are some mathematical foundations for them (Xu et al., ), standard accuracy-based probing has been criticized for combining what the probe can learn with what the representation actually encodes, making the results sensitive to probe capacity, data size, and hyperparameters (Hewitt and Liang, 2019; Pimentel et al., 2020). Further, high probe scores often come from exploiting superficial linear context cues rather than genuine structural knowledge (Kunz and Kuhlmann, 2020): probing can even reveal features that a model does not use for its task (Ravichander et al., 2021; Kumar et al., 2022). Consequently, some critiques motivate adopting information-theoretic lenses (Voita and Titov, 2020; Diego-Simón et al., 2025). Similar to some of these information-theoretic approaches (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017; Skean et al., 2025; Voita et al., 2019), our method makes use of information theory. For a more detailed discussion of related work, please refer to App. C.
@@ 2 Method
Given a prefix sequence of length $T$, ${\mathbf{x}}_{1:T}=\left[x_{1},x_{2},\dots,x_{T}\right]$, the decoder-only LM, $\mathcal{M}$, generates subsequent tokens $\{\widehat{x}_{T+\tau}\mid\tau=1,2,\dots\}$ autoregressively. We denote the hidden activation at position $t$ after the $\ell^{\text{th}}$ transformer layer by ${h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}}\in\mathbb{R}^{d}$ where $d$ is the hidden dimension of the model.
In our analysis, we treat the outputs $h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}$} of individual transformer blocks as the fundamental units of computation [^1] We use the terms “hidden state” and “(hidden) activation” interchangeably to refer to these outputs.. We posit that a block’s output is informative about the computation performed within that block (Pal et al., 2023; Shai et al., 2024; Li et al., 2023; Elhage et al., 2021; Lindsey et al., 2025). We use them as a proxy for the model’s internal processing since each block’s computation contributes to the overall function of the model. To analyze how planning emerges in the model’s internal dynamics, we examine collections of layer–token indices $\mathcal{S}$ and aggregate the corresponding block outputs as $G_{\mathcal{S}}=\{{h_{\scriptscriptstyle t}^{\scriptscriptstyle\ell}}\mid(\ell,t)\in\mathcal{S}\}$, a representation for the computations within its block. While $G_{\mathcal{S}}$ provides a fine-grained representation of block computations, it often contains unnecessary details for our planning analysis. Hidden states may encode small variations due to token positions, normalization shifts, or attention noise, factors that are critical for exact output reconstruction but irrelevant for studying planning structure. Moreover, MI estimates between such high-dimensional vectors have high variance and are hard to find. To address both issues, we construct coarse, discrete representations of $G_{\mathcal{S}}$. These abstractions preserve the structural dependencies required for our analysis while filtering out irrelevant micro-level variability, yielding more stable and interpretable MI estimates. We depict the proposed method in Fig. 1. In the first step, we train a model that maps each collection $G_{\mathcal{S}}$ to a coarse discrete representation $Z_{\mathcal{S}}$. Using the trained model, we then obtain dataset-wide codes for the target hidden states and use these to estimate mutual information between the coarse representations of LM’s internal computations.
Figure 1: The proposed method. Step 1 (training): For a frozen LM $\mathcal{M}$, hidden states from selected transformer blocks $G_{S}$ are passed through a VQ-VAE encoder, which maps each block to a latent vector and then to a discrete codebook index $Z_{S}\in[K]$, providing coarse summaries of internal computations. Step 2 (analysis): For two sets of hidden-state blocks, $G_{S_{A}}$ and $G_{S_{B}}$, we apply the trained encoder and codebook to the dataset to obtain discrete variables $Z_{A}$ and $Z_{B}$ and collect their empirical co-occurrence counts. Step 3: Using these statistics, we estimate joint and marginal distributions $p(z_{a},z_{b})$, $p(z_{a})$, and $p(z_{b})$ to compute mutual information $I(Z_{A};Z_{B})$ (and its normalized variant). Then, we analyze how information is shared between different components of the model’s computation.
@@
To investigate how planning is realized within the model, we use MI as a principled way to quantify relationships between the coarse representations of different computations, corresponding to Steps 2 and 3 of Fig. 1. Suppose we focus on two hidden state blocks, $G_{{A}}$ and $G_{{B}}$, which are random variables depending on the data distribution and the trained model parameters. We denote their coarse, discrete representations by $Z_{A}\in\mathcal{Z_{A}}$ and $Z_{B}\in\mathcal{Z_{B}}$ where $\mathcal{Z_{A}}$ and $\mathcal{Z_{B}}$ are the finite alphabets of possible codes. For realizations $z_{a}\in\mathcal{Z}_{A}$ and $z_{b}\in\mathcal{Z}_{B}$, $p(z_{a},z_{b})$, $p(z_{a})$, and $p(z_{b})$ refer to the joint probability of $Z_{A}$ and $Z_{B}$, and the marginal probabilities of $Z_{A}$ and $Z_{B}$, respectively. Mutual information, which measures how much knowing $Z_{A}$ reduces uncertainty about $Z_{B}$ (and vice versa), is defined as,
@@
Since the true data distribution is inaccessible, we approximate by replacing $p(z_{a})$, $p(z_{b})$, and $p(z_{a},z_{b})$ in Eq. 1 with their empirical estimates from the datasets, obtaining by counting code pairs as in Step 2 of Fig. 1. We analyze the finite-sample error of this count-based MI estimator (and a bias-corrected variant) in App. E. We do not interpret MI as a direct measure of planning; rather, we use MI comparatively. We further define normalized mutual information (nMI) metric as
@@
where $\mathcal{I}_{\max}$ denotes the maximum value across the set of MI calculations performed in the same experimental analysis for comparison. For our purpose of LM planning investigation, the nMI metric is sufficient since our analyses depend on relative, rather than absolute, comparisons between sets of computations. Accordingly, all claims are about within-setting trends, not about nMI as an absolute planning score. Absolute values can be misleading for three reasons: (i) the estimated MI of coarse representations is bounded above by the logarithm of the alphabet size, making it inherently dependent on the experimental choice; (ii) any coarse-graining necessarily discards some information, which is advantageous for our setting as discussed but removes any meaning of absolute MI values related to the original hidden states; and (iii) answering our three research questions (§ 3) requires only relative comparisons of MI instead of relying on absolute values.